# Consumption Partial Insurance

### v5

@@ 1 Introduction
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling than a positive shock, because it is farther away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and kurtosis suggest that for most households the realizations of income shocks tend to be either small (i.e., close to the zero mean) or far out in the left tail. However, these features vary by age and by the household’s position in the income distribution.
Second, in estimating BPP using the updated PSID in recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al. (2023) find a similarly low pass-through of permanent shocks. We discuss this further when we present our empirical results. We attribute about a third of the gap from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the gap arises because the modern PSID provides data biennially, versus annually previously, which tends to mute the transmission of higher frequency shocks. The last third reflects the different calendar times.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of large negative shocks versus positive or smaller ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. 2.) The pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie farther out in the left tail), but they also transmit more strongly as their magnitude worsens. 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
@@
With left skewness and excess kurtosis, large negative shocks are more likely and more severe than positive ones, so households would rather remove such tail risk from their portfolio. A back-of-the-envelop calculation suggests that households facing large downwards risk would give up 4%-8% of lifetime consumption when income risk transmits linearly and rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission. There are several estimates of the welfare costs of income risk in the literature. Two settings in which the baseline environment concerns idiosyncratic tail income risk of similar nature and severity to ours are Busch and Ludwig (2024) and Guvenen et al. (2024). Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk (a more distant counterfactual from ours). [^3] These papers calculate welfare costs using full structural models in which the consumption responses depend on endogenously accumulated wealth. By contrast, we rely on approximate consumption rules, which take wealth as given. We show in appendix D that this is unlikely to be quantitatively important., [^4] Welfare cost estimates of aggregate risk include Barro (2009), who find that agents can give up about 20% of GDP each year to eliminate rare disasters, and Georgarakos et al. (2025), who find that households can sacrifice 5%-6% of lifetime consumption to eliminate business cycle fluctuations.
@@
The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models. [^6] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand. Early works focus on self-insurance. Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance. More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021). Except Madera (2019), these papers abstract from higher income moments due to computational difficulties. De Nardi et al. (2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk increases partial insurance due to stronger precautionary motives. In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis; they distinguish between good and bad shocks and show that the latter are worse insured. Guvenen et al. (2024) find that tail shocks are worse insured than average shocks but the benchmark method in BPP overstates the true extent of insurance. Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; we measure aggregate insurance from all mechanisms together, while accounting for higher-order features of the income distribution, keeping our model parametrically unspecified.
@@
The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects. Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth, while Guvenen et al. (2021) provide even richer evidence from administrative data. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Backus et al., 2011; Ghosh and Julliard, 2012; Constantinides and Ghosh, 2017), business cycles (McKay, 2017), and labor market dynamics (Ai and Bhandari, 2021).
In the rest of the paper, section 2 presents the income and consumption processes and section 3 discusses identification. Section 4 discusses the implementation, section 5 presents the results, and section 6 concludes. The appendix has additional derivations and results.
@@ 2.1 Income process
where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race) and $\boldsymbol{\delta}$ is its loading factor. The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$. We assume that permanent income has a unit root, $P_{it}=P_{it-1}+\zeta_{it}$, where $\zeta_{it}$ is the permanent income shock. These modeling choices give rise to the familiar permanent-transitory formulation
$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vspace{-1ex}$ (3)
where $\Delta y_{it}=\Delta\ln Y_{it}-\Delta\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}$ is income growth net of the deterministic profile, i.e., unexplained income growth, and $\Delta$ is the first difference operator. [^8] See Meghir and Pistaferri (2011) for a comprehensive review of modern income processes.
@@
${\operatorname*{\mathbb{E}}}(\zeta_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{\zeta_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{\zeta_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{\zeta_{t}}&\text{for }m=4\\[-2.0pt] \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}(v_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{v_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{v_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{v_{t}}&\text{for }m=4.\\[-2.0pt] \end{cases}\vspace{-1ex}$
$\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction. $\gamma$ denotes the skewness/asymmetry of the distribution, i.e., the relative length of the upper and lower tails. Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
@@
As discussed in the introduction, several recent studies document that the cross-sectional distribution of income growth exhibits left skewness and excess kurtosis, i.e., a long and fat left tail. Our specification of the moments of income shocks enables our income process to be consistent with these facts, allowing us to subsequently decompose its second- and higher-order moments into the relative contributions of permanent and transitory shocks.
@@ 2.2 Linear consumption function
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. [^9] In addition, households in Blundell et al. (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds, the better they can self-insure. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. These include the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^13] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We first restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the income distribution.
@@
As we describe in section 4, the PSID provides data only every two years within our time period. This necessitates to recast the income and consumption growth processes of section 2 to this lower frequency, a straightforward operation that we carry out in appendix B. Nevertheless, for consistency with the earlier literature, we provide the identifying statements subsequently as if the data come at a yearly frequency. The complete identifying statements at the lower frequency of the modern PSID appear in appendix B.
@@ 3.1 Income process parameters
The second moment depends on the variance of permanent shocks at $t$ and the sum of variances of transitory shocks at $t$ and $t-1$; it thus reflects the dispersion of one permanent and two transitory shocks. The fourth moment reflects, similarly, the dispersion and kurtosis of both types of shocks. By contrast, the third moment is driven by skewness in the permanent shock. Barring strong non stationarities in the distribution of transitory shocks, the difference between $\gamma_{v_{t}}$ and $\gamma_{v_{t-1}}$ will be small (zero if the distribution is stationary), leaving $\gamma_{\zeta_{t}}$, the skewness in permanent shocks, to drive entirely the third moment of income.
@@ 3.2 Linear consumption function parameters
BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B.2. Formally, the transmission parameter of permanent shocks is identified as
@@ 3.3 Quadratic consumption function parameters
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^16] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^17] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^18] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
@@
$-{\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$

### v6

@@ 1 Introduction
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at 0.83-0.83 and 1.43-1.43, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling than a positive shock, because it is farther away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and kurtosis suggest that for most households the realizations of income shocks tend to be either small (i.e., close to the zero mean) or far out in the left tail. However, these features vary by age and by the household’s position in the income distribution.
Second, in estimating BPP using the updated PSID in recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al. (2023) find a similarly low pass-through of persistent shocks. We discuss this further when we present our empirical results. We attribute about a third of the gap from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the gap arises because the modern PSID provides data biennially, versus annually previously, which tends to mute the transmission of higher frequency shocks. The last third reflects the different calendar times.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of large negative shocks versus positive or smaller ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. 2.) The pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie farther out in the left tail), but they also transmit more strongly as their magnitude worsens. 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). These and previous findings are robust to allowing for shocks that have linearly persistent, though not necessarily permanent, effects on income. Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
@@
With left skewness and excess kurtosis, large negative shocks are more likely and more severe than positive ones, so households would rather remove such tail risk from their portfolio. A back-of-the-envelop calculation suggests that households facing large downwards risk would give up 4%-8% of lifetime consumption when income risk transmits linearly and rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission. There are several estimates of the welfare costs of income risk in the literature. Two settings in which the baseline environment concerns idiosyncratic tail income risk of similar nature and severity to ours are Busch and Ludwig (2024) and Guvenen et al. (2024). Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk. [^3] These papers calculate welfare costs using full structural models in which the consumption responses depend on endogenously accumulated wealth. By contrast, we rely on approximate consumption rules, which take wealth as given. We show in appendix D that this is unlikely to be quantitatively important., [^4] Welfare cost estimates of aggregate risk include Barro (2009), who find that agents can give up about 20% of GDP each year to eliminate rare disasters, and Georgarakos et al. (2025), who find that households can sacrifice 5%-6% of lifetime consumption to eliminate business cycle fluctuations.
@@
The second branch of the literature studies specific insurance mechanisms using calibrated consumption lifecycle models. [^6] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand. Early works focus on self-insurance: Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more. Recent studies examine the insurance implications of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), and added workers (Wu and Krueger, 2021). Except for Madera (2019), these papers abstract from higher income moments due to computational reasons. De Nardi et al. (2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk strengthens precautionary motives and increases partial insurance. In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis, distinguishing good from bad shocks and showing the latter are worse insured. Guvenen et al. (2024) find that tail shocks are worse insured than average ones but BPP’s benchmark method overstates insurance against them. Unlike this quantitative literature, we take no stance on specific insurance mechanisms; we measure aggregate insurance across all mechanisms while accounting for higher-order moments and keeping our model parametrically unspecified.
@@
The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects. Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth, while Guvenen et al. (2021) provide richer evidence from administrative data. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications for consumption. We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Backus et al., 2011; Ghosh and Julliard, 2012; Constantinides and Ghosh, 2017), business cycles (McKay, 2017), and labor market dynamics (Ai and Bhandari, 2021).
In the rest of the paper, section 2 introduces the income and consumption processes and section 3 discusses their identification. Section 4 presents the empirical implementation, section 5 discusses the results, section 6 extends the main analysis to a more general income process, and section 7 concludes. The appendix has additional derivations and results.
@@ 2.1 Income process
where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race) and $\boldsymbol{\delta}$ is its loading factor. The remaining terms make up the stochastic component of income, consisting of persistent income $P_{it}$ and transitory shock $v_{it}$.
We assume that $P_{it}$ has a unit root, namely $P_{it}=P_{it-1}+\zeta_{it}$, so in fact $P_{it}$ is permanent income and $\zeta_{it}$ is the permanent income shock. This modeling choice gives rise to the familiar permanent-transitory formulation
$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vskip-4.30554pt$ (3)
where $\Delta y_{it}=\Delta\ln Y_{it}-\Delta\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}$ is income growth net of the deterministic profile, i.e., unexplained income growth, and $\Delta$ is the first difference operator. [^8] See Meghir and Pistaferri (2011) for a comprehensive review of modern income processes. Section 6 verifies our main results using a more general income process that relaxes the unit root assumption.
@@
[…]
$\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction. $\gamma$ denotes the skewness/asymmetry of the distribution, i.e., the relative length of the upper and lower tails. Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $\sigma,\sigma]$, and thus characterizes the thickness of the tails.
@@
As discussed in the introduction, several recent studies document that the cross-sectional distribution of income growth exhibits left skewness and excess kurtosis, i.e., a long and fat left tail. Our specification of the moments of income shocks enables our income process to be consistent with these facts, allowing us to subsequently decompose the various income moments into the relative contributions of permanent and transitory shocks.
@@ 2.2 Linear consumption function
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. [^9] Households in Blundell et al. (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds, the better they can self-insure. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 3 Identification
[…]
@@
As we describe in section 4, the PSID provides data only every two years within our time period. This necessitates to recast the income and consumption growth processes of section 2 to this lower frequency, a straightforward operation that we carry out in appendices A and B. Nevertheless, for consistency with the earlier literature, we provide the identifying statements subsequently as if the data come at a yearly frequency. The complete identifying statements at the lower frequency of the modern PSID appear in appendix B.
@@ 3.1 Income process parameters
The second moment depends on the variance of permanent shocks at $t$ and the sum of variances of transitory shocks at $t$ and t1t-1; it thus reflects the dispersion of one permanent and two transitory shocks. The fourth moment reflects, similarly, the dispersion and kurtosis of both types of shocks. By contrast, the third moment is driven by skewness in the permanent shock. Barring strong non stationarities in the distribution of transitory shocks, the difference between $\gamma_{v_{t}}$ and $\gamma_{v_{t-1}}$ will be small (zero if the distribution is stationary), leaving $\gamma_{\zeta_{t}}$, the skewness in permanent shocks, to drive the third moment of income.
@@ 3.2 Linear consumption function parameters
BPP discuss identification in the linear function at length; we provide a summary here while we relegate the details to appendix B.2. Formally, the transmission parameter of permanent shocks is identified as
@@ 3.3 Quadratic consumption function parameters
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^16] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^17] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank –and hence invertible– for any distribution of income shocks other than Bernoulli (for which there is no empirical evidence anyway). [^18] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
@@
${\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$